The fraction (5x-11) / (2x 2 + x - 6) was obtained by adding the two fractions A / (x + 2) and B / (2x - 3) . The values of A and B must be, respectively: (a) 5x, -11, (b) -11, 5x, (c) -1, 3, (d) 3, -1, (e) 5, -11

The fraction (5x-11)/(2x2 + x - 6) was obtained by adding the two fractions A/(x + 2) and B/(2x - 3). The values of A and B must be, respectively:

(a) 5x, -11, (b) -11, 5x, (c) -1, 3, (d) 3, -1, (e) 5, -11

Grade:8

1 Answers

Nandana
7 years ago
option (d) is currect
Sol : –
A(2x-3) + B(x+2) = 2Ax-3A+Bx+2B /(2X^2+4X-3X-6)
= x(2A+B)-(3A-2B) /( 2X^2+4X-3X-6)
equate right hand side & left hand sides , then denominator is same &
equilate the numerator  coefficients , we get
2A+B = 5
3A-2B = 11
by solving these equations we get A=3 & B=-1